Example II Given Change in Yield on Notes with different coupons, but same maturity same example but use a 10% coupon insteadCreate a new table but this time with a 10% rate
Answer #1
Year Coupon rate 10% Initial yield 5% 6% yield
1 100               95.24               94.34
2 100               90.70               89.00
3 100               86.38               83.96
4 100               82.27               79.21
5 1100            861.88            821.98
Sum         1,216.47         1,168.49
Percentage change -3.94%
(1,168.49 - 1,216.47) ÷ 1,216.47 × 100

Work:

Year Cash flow PVF @ 5% Discounted cash flow
1 100 0.952380952 95.24
2 100 0.907029478 90.7
3 100 0.863837599 86.38
4 100 0.822702475 82.27
5 1100 0.783526166 861.88
Price of bond @ YTM 5% 1216.47
Year Cash flow PVF @ 6% Discounted cash flow
1 100 0.943396226 94.34
2 100 0.88999644 89
3 100 0.839619283 83.96
4 100 0.792093663 79.21
5 1100 0.747258173 821.98
Price of bond @ YTM 6% 1168.49
  • Price of a bond is the present value of all future cash flows from it discounted at a required rate of
    return. This required rate of return is the YTM.
  • YTM is the yield the bond-holder earns provided he holds the bond till maturity.
  • Cash flows from a bond are the periodic coupons and the maturity value on redemption.
  • In this case, coupon = 1000 × 10% = 100 (Bond coupon rate is always applied on the par value of the bond)
  • Cash flow in Year 5 = coupon + redemption value = 100 + 1000 = 1100